OpenStudy (anonymous):

write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form ax^2+bx+c=0. Only solution is 10

OpenStudy (anonymous):

the equation is blank_____=0

mathslover (mathslover):

Hi the equation having two roots a and b is represented as : (x-a)(x-b) but in this situation you have a=b=10 right ? replace a and b with 10 what do you get?

mathslover (mathslover):

(x-10)(x-10)=x^2+100-2(x)(10)=x^2+100-20x this is the equation